consider the vector A=sin(π÷2)i - sin(π÷2x)j, what is the divergence of the vector A


1
Expert's answer
2021-05-17T12:29:34-0400

Gives

"A=sin\frac{\pi}{2} \hat{i}+sin\frac{\pi}{2x}\hat{j}"

Where"sin\frac{\pi}{2}=1"

"\nabla. A=(\frac{d}{dx}\hat{i}+\frac{d}{dy}\hat{j}).(sin\frac{\pi}{2} \hat{i}+sin\frac{\pi}{2x}\hat{j})"

"\nabla. A=\frac{d}{dx}(1)+\frac{d}{dy}(sin\frac{\pi}{2x})""\nabla.A=0+0=0"

"\nabla .A=0"